4 years ago man was six times as old as his son but in five years he will be only three times as old as his son. What are their present ages
Answer by rothauserc(4718)

let m be the age of the man and s be the age of his son, then
:
1) m - 4 = 6 * (s - 4)
:
2) m + 5 = 3 * (s + 5)
:
solve both equations for m and set them equal to each other
:
6 * (s - 4) + 4 = 3 * (s + 5) - 5
:
6s - 24 + 4 = 3s + 15 - 5
:
6s - 20 = 3s + 10
:
3s = 30
:
s = 10
:
substitute for s in equation 2 (Note we could have chosen 1 also)
:
m + 5 = 3 * (10 + 5)
:
m = 40
:
*****************************************
The man is 40 and his son is 10
:
check the answers
:
40 - 4 = 6 * (10 - 4)
:
36 = 36
:
40 + 5 = 3 * (10 + 5)
:
45 = 45
:
answers check
: